On December 31, 2020, after the adjusting entries were posted, the Accumulated Depreciation—Machinery account of Klopp had a balance of $882,000. During 2020, one of the machines that cost $387,000 was sold for $215,000 cash. This resulted in a loss of $20,000. Depreciation expense for 2020 was $72,000. Assuming that no other assets were disposed of during the year, how much was the accumulated depreciation balance on January 1, 2020?

Multiple Choice

  • $942,000

  • $962,000

  • $822,000

  • $802,000
Answer #1

Answer

  • Correct Answer = Option #2: $ 962,000
  • Working

Step #1

A

Sale Price of machine

$215,000

B

Loss on sale

$20,000

C = A+B

Book Value of machine sold

$235,000

D

Cost of machine sold

$387,000

E = D - C

Depreciation balance of machine sold

$152,000

Step #2

A

Accumulated Depreciation - 31 Dec 2020

$882,000

B

Depreciation balance of machine sold [calculated in Step #1]

$152,000

C

Depreciation expense for 2020

$72,000

D = A+B - C

Accumulated Depreciation - 1 Jan 2020

$962,000 [Answer]